The data below represents the overall miles per gallon​ (MPG) of 2008 SUVs priced under​ $30,000. Complete parts a and b below. Start 2 By 1 Matrix 1st Row 1st Column 21 comma 20 comma 22 comma 23 comma 18 comma 19 comma 18 comma 16 comma 18 comma 20 comma 20 comma 2nd Row 1st Column 16 comma 22 comma 19 comma 16 comma 17 comma 16 comma 18 comma 16 comma 20 comma 15 comma 22 EndMatrix a. Construct a 99​% confidence interval estimate for the population mean miles per gallon of 2008 SUVs priced under​ $30,000, assuming a normal distribution.
Answer #1

Solution:-

from the Given'data,

Mean(x̅)=18.73

Standard Deviation(s)=2.374 (use excel or calculator for the calculation)

Significance level(α)=0.01

n=22

Degree of freedom=22-1=21

then, tα/2,21=2.83136

= 13-73 ± 2.83 136 2.374 - 18.73+ I-4667& -07. 2632, 20-1968) 2/